OpenStudy (anonymous):

A machine starts dumping sand at the rate of 20 m3/min, forming a pile in the shape of a cone. The height of the pile is always twice the length of the base diameter. A. After 5 minutes, what is the height of the pile? B. After 5 minutes, how fast is the height increasing? C. After 5 minutes, how fast is the base radius increasing? D. After 10 minutes, how fast is the area of the base increasing?

OpenStudy (anonymous):

help anyone?

OpenStudy (cwrw238):

do you know the formula for the volume of a cone ?

OpenStudy (anonymous):

@cwrw238 i dont know how to solve for the rate because it does not give a height.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

V=1/3pir^2h

OpenStudy (anonymous):

this is what i have so far... h=2d h=4r r=h/4 dV/dt=20mcu./min 20=pi/48(h^3)d/dt=pi/48(3h^2)dh/dt

OpenStudy (anonymous):

this is where i got stuck!

OpenStudy (cwrw238):

V = pi r^2 h ------- 3 h = 2 * diameter of the base = 4 * radius so r = h/4 so we can write V = ( pi *( h/4)^2 * h ) / 3 V = pi h^3 ----- 48

OpenStudy (anonymous):

you forgot to take the derivative of h^3, no?

OpenStudy (cwrw238):

i dont need it to do A

OpenStudy (anonymous):

okay, then i have done that [art already. but this is where i got stuck!

OpenStudy (anonymous):

do we multiply 5 by 20, then solve for h? using V = pi h^3 ----- 48

OpenStudy (anonymous):

@cwrw238

OpenStudy (cwrw238):

yes 100 = pi* h^3 ------- 48

OpenStudy (anonymous):

i got 11.52, which doesnt make sense.

OpenStudy (anonymous):

@cwrw238 are you there?

OpenStudy (anonymous):

@mathstudent55 will youhelp me too?

OpenStudy (cwrw238):

hmmm does seem a high number - can you see anything wrong with my algebra?

OpenStudy (anonymous):

no. but i need to finish this. so i'll just go with that. can we go on to B, now please?

OpenStudy (cwrw238):

ok dV/dt = dV/dh * dh / dt we need to find dh/dt so dV/dt = 20 and dV/dh = 3 pi h^2 ------- = pi h^2/ 16 = 11.52^2 pi / 16 48 so plug these 2 values into the above formula to get dh/dt

OpenStudy (cwrw238):

agreed?

OpenStudy (anonymous):

agreed!

OpenStudy (anonymous):

wait! this part is confusing: 11.52^2 pi / 16

OpenStudy (anonymous):

why are we doing that???

OpenStudy (anonymous):

@cwrw238

OpenStudy (cwrw238):

h= 11.52 at time 5 minutes

OpenStudy (anonymous):

but how does that solve for dh/dt

OpenStudy (cwrw238):

we need to find dV/dh

OpenStudy (anonymous):

i got 26.05

OpenStudy (cwrw238):

20 = [11.52^2 pi / 16] * dh/dt dh/dt = 20* 16 ----- 11.52^2 pi

OpenStudy (anonymous):

.768 ???

OpenStudy (cwrw238):

i get that 0.768 m / min yes

OpenStudy (anonymous):

ok! C now?

OpenStudy (anonymous):

we need to find that derivative of the radius? right?

OpenStudy (cwrw238):

you do that in a similar way to the height dV/dt = dV/dr * dr / dt

OpenStudy (cwrw238):

find V in terms of r then differentiate

OpenStudy (cwrw238):

h = 4r so V = pi r^2 * 4r -------- 3 so what dv/dr ?

OpenStudy (anonymous):

dv/dr= 8rpi/3 ???

OpenStudy (cwrw238):

dV / dr = 4 pi r^2 ------ right? 3

OpenStudy (anonymous):

do you take the derivative of r^2 ??

OpenStudy (cwrw238):

no you take the derivative of (4 pi r^3) / 3

OpenStudy (anonymous):

oh. ok. i get it. i forgot about the r^3

OpenStudy (cwrw238):

sorry i made a mistake dV / dr = 4 pi r^2

OpenStudy (cwrw238):

its 4 * 3 pi r^(3-1) --- 3

OpenStudy (anonymous):

i already did that, haha.

OpenStudy (cwrw238):

good for you so we have 20 = 4pi r^2 * dr/dt dr/dt = 20 / 4 pi r^2 so we sub in r what is the value of r?

OpenStudy (anonymous):

the heigh over 4 so::: 11.52/4

OpenStudy (anonymous):

@cwrw238 right?

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

dr/dt=.012m/min

OpenStudy (anonymous):

@cwrw238 im gonna try to do D by myself. if need help i'll ask okay?

OpenStudy (anonymous):

thanks a bunch to you! rates of change are fun! so thank for helping!!!!

OpenStudy (cwrw238):

sorry gotta go now

OpenStudy (cwrw238):

dr/dt = 0.192 m / sec

OpenStudy (cwrw238):

in fact there was no need to do all that calculus if dh/dt = 0.768 then dr/dt = 0.768/4 = 0.192 !!

OpenStudy (cwrw238):

as for the area the rate of increase will be pi r^2 m^2/min that is pi * 0.192^2 = 0.116 cm^2 / min